OpenStudy (anonymous):

Can you help me with this chart!? Start with 40 candies on the reactant side of the paper and zero candies on the product side. Each round, you will be exchanging candies between R and P. For each round, half of the candies that were on the reactant (R) side of the paper from the last round should be moved to the product side. One-fourth of the candies that were on the product (P) side of the paper from the last round should be moved to the reactant side. (If you end up with a decimal for the number to exchange, round up.)

OpenStudy (aaronq):

R P 40 (move 20) 0 (move 0) 20 (move 10) 20 (move 5) 10+5=15 (move 7.5) 15+10=25 (move 6.25) 7.5+6.25=13.75 (move 6.875) etc